OpenStudy (anonymous):

WILL MEDAL; Given that the first 5 terms of a geometric sequence are 3, x, 12, y, and 48, find, x and y. Assume both x and y are positive.

OpenStudy (triciaal):

for a geometric sequence the terms have a common ratio example term 3/term 2 = term 5/ term 4

OpenStudy (triciaal):

set up 2 sets of ratio and solve the simultaneous equation

OpenStudy (cwrw238):

the 5 th term is 48 so a1* r^4 = 48 where a1 = 3 from this you can find the common ratio r

OpenStudy (triciaal):

* not simultaneous equation just quadratic and use the positive value

OpenStudy (triciaal):

I have x = 6 and y = 7 the ratio r = 2

OpenStudy (anonymous):

the second term would be 4?

OpenStudy (triciaal):

no your 2nd term is x

OpenStudy (triciaal):

first 5 terms of a geometric sequence are 3, x, 12, y, and 48

OpenStudy (anonymous):

im not quite sure on how to figure it out @triciaal

OpenStudy (triciaal):

48/y = y/12 = y/12 = 12/x = x/3 do you see this ?

OpenStudy (anonymous):

alright, would it be 6?

OpenStudy (triciaal):